OpenStudy (anonymous):

the second angle in a triangle is two third as large as the first. the third angle is one half as large as the second angle. find the angle measures.

OpenStudy (anonymous):

Let's call them A (1st) B(2nd) and C(3rd) B = 2/3 A C = 1/2 B Substitute C = 1/2 ( 2/3 A) C = 1/3A So: A = A B = 2/3A C = 1/3A That makes a total of 2A Total = 180 deg 2A = 180 A=90 B = 2/3 * 90 = 60 C = 1/3 * 90 = 30